OpenStudy (anonymous):

Let f(x)=12/4x+2 find f(-1) A. -6 B. -2 C. 2 D. 6

OpenStudy (princeharryyy):

f(-1) = 12/(4*(-1) + 2) => 12/(-2) = -6 @Morgan27

OpenStudy (joftheworld):

plug in (-1) into 4(x)+2 then divide that by 12

OpenStudy (joftheworld):

and if you have more problems like that just do the same

OpenStudy (anonymous):

Thank you. I got -6 at first but the way I did it just didn't seem right lol

OpenStudy (tkhunny):

AS WRITTEN f(-1) = 12/4(-1)+2 = 3(-1)+2 = -3+2 = -1 If you intended f(x) = 12/(4x+2), we have f(-1) = 12/(4(-1)+2) = 12/(-4+2) = 12/(-2) = -6 The parentheses are NOT optional if you want to write what you mean.

OpenStudy (anonymous):

Thank you for explaining it. :)

OpenStudy (anonymous):

Could you help me with one more? Pleaseee :)

OpenStudy (tkhunny):

Post new problems on new threads, please. SHOW your work. Don't make us guess.